OpenStudy (bloomlocke367):

In March, a family starts saving for a vacation they are planning for the end of August. The family expects the vacation to cost $1375. The start with $125. Each month they plan to deposit 20% more than the previous month. Will they have enough money for their trip? If not, how much more do they need?

OpenStudy (bloomlocke367):

@iGreen

OpenStudy (blake57roger):

they have enough money for their trip

OpenStudy (bloomlocke367):

Please don't give me a direct answer... that doesn't help, and it's against the rules.

OpenStudy (anonymous):

Do you know the process for solving this problem?

OpenStudy (anonymous):

125 + 20% * 5

OpenStudy (bloomlocke367):

I'm pretty sure I need to use the summation notation of some sort

OpenStudy (bloomlocke367):

and wouldn't it be 120%? 20% would decrease the value... that makes no sense.

OpenStudy (bloomlocke367):

Don't I need to use \[S _{n}=\frac{ a _{1}(1-r ^{n}) }{ 1-r }\]

OpenStudy (bloomlocke367):

@CountryGurl15

OpenStudy (anonymous):

Um, yes you are on the right track =) I thought your problem said 20% tho =)

OpenStudy (bloomlocke367):

So would \[a_1=125\] \[n=5\] and \[r=1.2\] ?

OpenStudy (anonymous):

Yes that looks right =)

OpenStudy (anonymous):

So solve now, and see what you get =)

OpenStudy (bloomlocke367):

\[S_5=\frac{ 125(1-1.2^5) }{ 1-1.2 }\]

OpenStudy (anonymous):

Yup, now u hafta just solve

OpenStudy (bloomlocke367):

Sorry! my openstudy was glitchy. so we have\[S_5=\frac{ 125(1-2.48832) }{1-1.2 }\] \[=\frac{ 125(-1.48832) }{ -0.2}\] \[=\frac{ -186.04 }{ -0.2 }\] \[=930.20\]

OpenStudy (bloomlocke367):

So they don't have enough money?..

OpenStudy (bloomlocke367):

@CountryGurl15

OpenStudy (anonymous):

Um, no they don't apparently =)

OpenStudy (bloomlocke367):

Okay.... is that correct?

OpenStudy (bloomlocke367):

Thanks for your help.

OpenStudy (anonymous):

Yes that is correct, and No problem =)